# Specific heat of Gd3+- and Eu2+-based magnetic compounds

### v1

@@ Abstract
We have studied theoretically the specific heat of a large number of non-frustrated magnetic structures described by the Heisenberg model for systems with total angular momentum $J=7/2$, which corresponds to the 4f7 configuration of Gd+3 and Eu+2. For a given critical temperature (controlled by the magnitude of the exchange interactions) we find that to a large accuracy, the specific heat is controlled by two parameters: the effective number of neighbors (which controls the importance of quantum fluctuations) and the axial anisotropy. Using these two parameters we fit the specific heat of four Gd compounds and two Eu compounds obtaining a remarkable agreement. Another Gd compound has been fitted previously. Our work opens the possibility of describing the specific heat of non-frustrated 4f7 systems within a general framework.
@@ I Introduction
Rare-earth-based materials are of great interest in both basic and applied condensed matter physics, showing a plethora of captivating fundamental physical phenomena, like unconventional superconductivity Mathur et al. (1998); Shioda et al. (2021), Kondo effect Kobayashi et al. (2008); Romero et al. (2013); Magnavita et al. (2016), quantum criticality Alvarez et al. (2004), quadrupolar order and frustration Watanuki et al. (2005); Okuyama et al. (2005); Ji et al. (2007); Song et al. (2020); Franco et al. (2024).
@@
However, an exception is expected for compounds containing Gd+3 and Eu+2, both of which correspond to the 4f7 configuration. In this case, the ground-state multiplet constructed by Hund rules is ${}^{8}S_{7/2}$, which correspond to total spin $S=7/2$, total orbital angular momentum $L=0$ and total angular momentum $J=7/2$. Since neither the spin nor terms with $L=0$ are affected by crystal field, one expects that the only source of splitting of the ground-state octuplet is the exchange interaction between rare-earth ions responsible for the magnetic structure of the compound. This is true as a first approximation. In fact we will show that the specific heat of five Gd compounds can be well fitted ignoring crystal-field effects.
Nevertheless, as a consequence of the spin-orbit coupling, the ground-state multiplet acquires some admixture of ${}^{6}P_{7/2}$ with $L=1$. The explicit calculation has been done in Appendix B of Ref. Betancourth et al., 2019, where it has been estimated that the amount of the excited multiplet in the ground state is $\sim 3\%$. Terms with $L=1$ are not affected by cubic crystal fields but are sensitive to lower symmetries. In particular, they are affected by an axial term, which (using Wigner-Eckart theorem) can be written in the form $H_{K}=K[3J^{2}_{z}-J(J+1)]$, where $K$ is expected to be small.
@@
Even if $H_{K}$ can be neglected, systematic deviations of the observed magnetic contribution to the specific heat $C(T)$ in Gd-based compounds from the corresponding mean field result were observed since long time ago, pointing to the importance of quantum fluctuations. The early measurements of $C(T)$ on GdNi5 Szewczyk et al. (1992) were compared with calculations in the molecular field approximation for $T<T_{C}=32$ K, where $T_{C}$ is the critical temperature, with good agreement. However the authors recognize evident discrepancies at low temperature and in the vicinity of $T_{C}$ that can be attributed to “short-range ordering effects not taken into account”. Simultaneous measurements of $C(T)$ on other Gd compounds Bouvier et al. (1991) show more significant deviations from the mean-field result. The origin of such deviations was investigated introducing modulations in the amplitude of the magnetic moments in the mean-field approximation Blanco et al. (1991). Such a model describes different possibilities of the observed temperature dependence of $C(T)$ for $T<T_{C}$. However, some compounds clearly escape from those predictions specially concerning observed magnetic fluctuations right above $T_{C}$.
@@
Recent research on Eu2+ based compounds show similar shortcomings of the mean-field approach Kumar et al. (2010); Giovannini et al. (2021). The fact that Eu2+ has a much larger atomic volume and therefore may participate in the formation of very different crystalline structures indicates that the observed deviations is an intrinsic characteristic of the magnetic interactions and not dependent from crystalline symmetries nor the sign of the interactions. Therefore, for a quantitative explanation of the observed $C(T)$, it is necessary to include quantum fluctuations and go beyond mean field.
@@
In this work we report the specific heat of a large number of structures described by a Heisenberg Hamiltonian containing exchange interactions at different distances
$H_{0}=\sum_{i,\delta}J_{\delta}\mathbf{S}_{i}\cdot\mathbf{S}_{i+\delta}/2,$ (1)
@@
where $\mathbf{S}_{i}$ denotes the spin 7/2 at site $i$, $\delta$ labels the non-equivalent space vectors connecting different spins at short distances (nearest and possibly further neighbors) and $J_{\delta}$ is the corresponding exchange interaction.
@@
The theoretically studied magnetic structures are non frustrated. They correspond to either ferromagnetic interactions or non-frustrated antiferromagnetic ones. We find that the specific heat of all these structures can be characterized to a large degree of accuracy by two parameters: i) the effective number of neighbors (to be defined more precisely below) and ii) the magnetic anisotropy defined by $H_{K}$ at each site. Using these two parameters we fit the reported specific heat as a function of temperature for six compounds obtaining an impressive agreement.
@@
The effect of fluctuations on the specific heat of a spin system described by the Heisenberg model is expected to decrease as the number of neighboring spins increases. We can draw an analogy to a binary random distribution (which would correspond to spin 1/2). It is known that after $N$ attempts in a binary random distribution $x$ with probabilities $p$ and $1-p$ for the values $x=1$ and $x=-1$ respectively, the mean value (proportional to the magnetic moment) is $\langle x\rangle=(2p-1)N$ and the standard deviation is $\sqrt{\langle x^{2}\rangle-\langle x\rangle^{2}}=2\sqrt{Np(1-p)}$. Therefore, the ratio between the standard deviation and the mean value decreases as $1/\sqrt{N}$. Similarly in our quantum Heisenberg model, one expects that as the number of neighbors increases, the specific heat is more similar to the mean-field result, in which quantum fluctuations are neglected and the effective magnetic field that each ion feels is replaced by its mean value.
@@
Generalizing the above results to the case of a distribution with intensities ±$\pm|J_{\delta}|$ for $z_{\delta}$ neighbors we obtain after some algebra (see Appendix)
@@
$\frac{\langle x^{2}\rangle-\langle x\rangle^{2}}{\langle x\rangle^{2}}=\frac{4p(1-p)}{2p-1}r\text{, where~{}~{}}r=\frac{\Sigma_{\delta}z_{\delta}J_{\delta}^{2}}{(\Sigma_{\delta}z_{\delta}|J_{\delta}|)^{2}}=\frac{1}{z}.$ (2)
@@
Here, $z_{\delta}$ represents the number of neighbors of type $\delta$ (with exchange interaction $J_{\delta}$), and the summation goes over all types of neighbors. The quantity $z=1/r$ defined by Eq. (2) captures the effective number of neighbors, taking into account the strength ($J_{\delta}^{2}$) and number ($z_{\delta}$) of each type of interaction. Remarkably, for unfrustrated systems, the observed specific heat depends mainly on this effective number of neighbors $z$, independent of the specific lattice structure. For larger $z$ (higher coordination number), the results approach the mean-field ones, where quantum fluctuations are suppressed. Conversely, for lower $z$ (fewer neighbors), quantum fluctuations become more important, leading to a smaller specific heat below the critical temperature $T_{C}$ and a larger specific heat above $T_{C}$. Explicit calculations are shown in Section III.
@@
Moreover, for some actual systems, the effect of the axial anisotropy $H_{K}$ becomes important, redistributing the entropy below $T_{C}$ (see Section IV).
@@
Using the above two parameters for simple lattices, and the known value of $T_{C}$, we have fitted the data for the specific heat of GdNi3Ga9, GdCoIn5, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. The corresponding results are presented in Section VI. Using essentially the same method, the specific heat of GdCoIn5 has been fitted in Ref. Facio et al., 2015.
@@ II Methods
We have calculated the specific heat of a large number of magnetic structures form the numerical derivative of the internal energy as a function of temperature. This energy was obtained using quantum Monte Carlo (QMC) simulations from the ALPS libraries (“dirloop_sse” package) Albuquerque et al. (2007); Bauer et al. (2011). Our calculations encompassed a wide range of system sizes, reaching up to $8\times 30^{3}$ magnetic moments.
To generate the lattice configurations for the ALPS simulations, we utilized the “SUNNY” package Dahlbom et al. (2024) in conjunction with crystallographic information files (CIF) obtained from public repositories. Specifically, CIF files for Fe in simple cubic (SC), body-centered cubic (BCC), hexagonal close-packed (HCP), and face-centered cubic (FCC) structures, as well as C in diamond, MgAl2Cu (for EuPdSn2), and ErNi3Al9, were retrieved from the Materials Project Jain et al. (2013) database. The ErNi3Al9 CIF file was obtained from the Crystallography Open Database Gražulis et al. (2009) using data from reference Gladyshevskii et al., 1993. In each case, the magnetic structure was created by replacing C, Fe, Mg, or Er with a spin moment $S=7/2$.
For all lattices, we adopted a rectangular cuboid supercell geometry as the conventional unit cell. This choice resulted in a varying number of sites within the supercell, with the diamond lattice requiring the largest (8 sites). The simulations were performed on systems containing $N^{3}$ supercells with periodic boundary conditions, typically with $N=10$, $20$, and $30$. For cases exhibiting size effects, the results were extrapolated to the thermodynamic limit assuming a $1/N$ dependence.
The well-known sign problem Pan and Meng (2024) associated with QMC simulations limited our investigation to non-frustrated magnetic structures. At low temperatures, the number of Monte Carlo steps required for statistically reliable results increases exponentially. Consequently, the internal energy was primarily computed in a temperature range between $0.2T_{C}$ and $2T_{C}$, where $T_{C}$ is the critical temperature of the system. In a few cases, the simulations were extended to $0.1T_{C}$ at low temperatures. However, the reliability of the results below $0.2T_{C}$ is limited.
@@
Our theoretical results for the specific heat for all lattices used here are available in Garcia, D. J., Sereni, J., Aligia, A. A., .
@@ III Role of the effective number of neighbors
In this Section we report our results for the specific heat and critical temperature for several simple lattices as a function of the number of neighbors that interact with a given site. For simplicity in all lattices considered, all sites are equivalent, and we assume all $J_{\delta}=J_{1}$, where $J_{1}$ is the NN interaction, up to a certain distance $d$ and 0 for longer distances. Therefore the effective number of neighbors $z=\Sigma_{\delta}z_{\delta}$ is just the sum of all neighbors within a sphere of radius $d$. In Section V we discuss more general cases showing that the results are not severely affected by this assumption.
@@
Most of our studies assume ferromagnetic interactions ($J_{1}<0$). In general, the results for the corresponding non-frustrated antiferromagnetic structures are very similar, particularly for $z>6$, except at very low temperatures, for which the specific heat dominated by magnons behaves as $T^{3/2}$ ($T^{3}$) for ferro- (antiferro-) magnetic structures.
@@
In Fig. 1 we show the critical temperature $T_{C}$ as a function of $z$ for the different structures considered. The calculations have been done in finite systems containing $N^{3}$ conventional unit cells, with $N=10$, 20 and 30, as explained in Section II. The extrapolated value is also indicated in the figure.
@@
As expected, $T_{C}$ increases almost linearly with $z$ as the mean-field value $T_{C}^{\text{ MF}}$ (the effective magnetic field due to the interactions is proportional to $zJ_{1}$, see Section IV.1), but it is lower than $T_{C}^{\text{ MF}}$, particularly for low $z$ due to the effect of the quantum fluctuations, which is larger for low $z$. For large enough $z$ one expects that $T_{C}$ and $T_{C}^{\text{ MF}}$ coincide. However, for large $z$ in a finite system with periodic boundary conditions, the effective coordination is smaller when the range of the interactions $d$ becomes of the order of the system size, because neighbors at different distances coincide, and it is not possible to reach this limit with our method.
@@
Figure 2: (Color online) Specific heat over temperature as a function of temperature for several ferromagnetic structures with different effective number of neighbors $z$. Top:$z\leq 12$. Bottom: $z\geq 12$. There is an almost perfect superposition between HCP and FCC for $z=12$, among SC, FCC and HCP for $z=18$, and among SC, BCC and the EuPdSn2 structure for $z=26$. Therefore only one of the curves for each of these three sets is shown. LHc(4) corresponds to the layered honeycomb structure with $z=4$.
@@
In Fig. 2 we show the results for the specific heat as a function of temperature for a large number of ferromagnetic structures with different number of magnetic neighbors $z$. Also shown is the mean-field result which corresponds to the limit $z\rightarrow\infty$. This result shows the larger $C$ for $T<T_{C}$, while $C=0$ for $T>T_{C}$. In contrast, $C$ for a honeycomb layered system with $z=4$ (three NN within the plane of the honeycomb layer and one above or below the layer in alternating sites) shows the smaller specific heat below $T_{C}$ and the largest one above $T_{C}$. Note that the diamond structure, which is completely different from the previous one, but also has $z=4$, has a very similar specific heat.
@@
Increasing $z$ from 4 to 6, the specific heat of the simple cubic structure with six NN displays the expected trend of increasing (decreasing) $C$ below (above) $T_{C}$. Increasing $z$ up to 14 using the bcc structure with 8 NN and 6 next NN (NNN), the same trend is in general observed. However, the specific heat of the honeycomb lattice (with the four links as described before and adding 6 second neighbors in the plane) with $z=10$ lies slightly below the cubic bcc with $z=8$ NN for $T\sim 0.5T_{C}$. In addition, somewhat unexpected is the result for the simple cubic structure including 8 NN and 6 third NN, whose $C$ at $T\sim 0.5T_{C}$ is greater than the corresponding one of the above mentioned BCC structure with the same $z=14$, and also greater than the specific heat of one structure with $z=16$ and two structures with $z=18$. Nevertheless, except for this case and the above mentioned inversion between $z=8$ and 10, the specific heat is very similar for different structures with the same $z$ and increases with increasing $z$ at intermediate temperature $T\sim 0.5T_{C}$.

### v2

@@ Abstract
We have studied theoretically the specific heat of a large number of non-frustrated magnetic structures described by the Heisenberg model for systems with total angular momentum $J=7/2$, corresponding to the 4f7 configuration of Gd+3 and Eu+2. For a given critical temperature (determined by the magnitude of the exchange interactions), we find that, to a high degree of accuracy, the specific heat is governed by two primary parameters: the effective number of neighbors, which dictates the extent of quantum fluctuations, and the axial anisotropy. Using these two parameters we fit the specific heat of four Gd compounds and two Eu compounds, achieving a remarkable agreement. Our work opens the possibility of describing the specific heat of 4f7 systems within a general framework.
@@ I Introduction
Rare-earth-based materials are of great interest in both basic and applied condensed matter physics, exhibiting a plethora of captivating fundamental physical phenomena, such as unconventional superconductivity[1, 2], Kondo effect [3, 4, 5], quantum criticality [6], quadrupolar order, and frustration[7, 8, 9, 10, 11].
@@
However, an exception is expected for compounds containing Gd+3 and Eu+2, both of which correspond to the 4f7 configuration. In this case, the ground-state multiplet constructed by Hund’s rules is ${}^{8}S_{7/2}$, which corresponds to total spin $S=7/2$, total orbital angular momentum $L=0$, and total angular momentum $J=7/2$. Since neither the spin nor terms with $L=0$ are affected by crystal fields, one expects that the only source of splitting of the ground-state octuplet is the exchange interaction between rare-earth ions responsible for the magnetic structure of the compound. This is true as a first approximation. In fact, we will show that the specific heat of five Gd compounds can be well fitted ignoring crystal-field effects.
Nevertheless, due to spin-orbit coupling, the ground-state multiplet acquires a small admixture of the ${}^{6}P_{7/2}$ state with $L=1$. An explicit calculation for Gd+3 presented in Appendix B of Ref. 12 estimates that the excited multiplet contributes approximately $3\%$ to the ground state. This admixture introduces a component with $L=1$ in the ground state, allowing the ground state to split under the influence of crystal fields if the symmetry at the rare-earth site is not high, with only a single axis of higher order. In such cases, the anisotropy term $H_{A}=C[3L^{2}_{z}-L(L+1)]$ becomes relevant. Note that in high-symmetry environments such as tetragonal, cubic, or octahedral point groups ($T$, $T_{d}$, $T_{h}$, $O$, $O_{h}$), the $L=1$ states remain unsplit. This result, derived from group theory, is intuitively understood by noting that a basis for $L=1$ states can be chosen to transform like the unit vectors $\hat{x}$, $\hat{y}$, $\hat{z}$, which are equivalent in high-symmetric environments. In contrast, if the maximum order axis of the point group is 2, a term $L^{2}_{x}-L^{2}_{y}$ is allowed. In even lower symmetries, $L_{x}L_{y}$ may also arise; however, their intensity is expected to be much smaller, and we can fit experimental data without including them.
Using the Wigner-Eckart theorem, $H_{A}$ can be written in the form $H_{A}=K[3J^{2}_{z}-J(J+1)]$, where $K$ is of order $\sim 3\%C$.
@@
Even if $H_{A}$ can be neglected, systematic deviations of the observed magnetic contribution to the specific heat $C(T)$ in Gd-based compounds from the corresponding mean-field result have been observed since long time ago, pointing to the importance of quantum fluctuations. Early measurements of $C(T)$ on GdNi5 [13] were compared with calculations in the molecular field approximation for $T<T_{C}=32$ K, where $T_{C}$ is the critical temperature, with good agreement. However, the authors recognized evident discrepancies at low temperature and in the vicinity of $T_{C}$ that can be attributed to “short-range ordering effects not taken into account”. Simultaneous measurements of $C(T)$ on other Gd compounds [14] show more significant deviations from the mean-field result. The origin of such deviations was investigated by introducing modulations in the amplitude of the magnetic moments in the mean-field approximation [15]. Such a model describes different possibilities of the observed temperature dependence of $C(T)$ for $T<T_{C}$. However, some compounds clearly escape from those predictions, especially concerning observed magnetic fluctuations right above $T_{C}$.
@@
Recent research on Eu2+-based compounds shows similar shortcomings of the mean-field approach [16, 17]. The fact that Eu2+ has a much larger atomic volume and therefore may participate in the formation of very different crystalline structures indicates that the observed deviations are an intrinsic characteristic of the magnetic interactions and not dependent on crystalline symmetries or the sign of the interactions. Therefore, for a quantitative explanation of the observed $C(T)$, it is necessary to include quantum fluctuations and go beyond mean field. The effect of these fluctuations was calculated recently using cluster mean-field theory [18].
@@
In this work, we report the specific heat of a large number of structures described by a Heisenberg Hamiltonian $H$ containing exchange interactions at different distances and an axial anisotropy term $H_{A}$:
$H=H_{0}+H_{A},$ (1)
where
$H_{0}=\sum_{i,\delta}J_{\delta}\mathbf{S}_{i}\cdot\mathbf{S}_{i+\delta}/2,\text{ and }H_{A}=K\sum_{i}[3S_{iz}^{2}-S(S+1)].$ (2)
@@
$\mathbf{S}_{i},S_{iz}$ denote the spin (or angular momentum) $|S|=7/2$ operators at site $i$ and its projection, $\delta$ labels the non-equivalent space vectors connecting different spins at short distances (nearest and possibly further neighbors), and $J_{\delta}$ is the corresponding exchange interaction. We use $S$ instead of $J$ to denote the operators to avoid confusion with the exchange interactions.
@@
The theoretically studied magnetic structures are non-frustrated and correspond to either ferromagnetic interactions or non-frustrated antiferromagnetic ones. We find that the specific heat of all these structures can be characterized to a large degree of accuracy by two parameters: i) the effective number of neighbors (to be defined more precisely below) and ii) the magnetic anisotropy defined by $H_{A}$. Using these two parameters, we fit the reported specific heat as a function of temperature for six compounds, obtaining an impressive agreement.
@@
The effect of fluctuations on the specific heat of a spin system described by the Heisenberg model is expected to decrease as the number of neighboring spins increases. We can draw an analogy to a binary random distribution (which would correspond to spin 1/2). It is known that after $N$ attempts in a binary random distribution $x$ with probabilities $p$ and $1-p$ for the values $x=1$ and $x=-1$, respectively, the mean value (proportional to the magnetic moment) is $\langle x\rangle=(2p-1)N$, and the standard deviation is $\sqrt{\langle x^{2}\rangle-\langle x\rangle^{2}}=2\sqrt{Np(1-p)}$. Therefore, the ratio between the standard deviation and the mean value decreases as $1/\sqrt{N}$. Similarly, in our quantum Heisenberg model, one expects that as the number of neighbors increases, the specific heat is more similar to the mean-field result, in which quantum fluctuations are neglected and the effective magnetic field that each ion feels is replaced by its mean value.
@@
Generalizing the above results to the case of a distribution with intensities ±$\pm|J_{\delta}|$ for $z_{\delta}$ neighbors, we obtain after some algebra (see Appendix):
@@
$\frac{\langle x^{2}\rangle-\langle x\rangle^{2}}{\langle x\rangle^{2}}=\frac{4p(1-p)}{(2p-1)^{2}}r,\quad\text{where}\quad r=\frac{\sum_{\delta}z_{\delta}J_{\delta}^{2}}{\left(\sum_{\delta}z_{\delta}|J_{\delta}|\right)^{2}}=\frac{1}{z},$ (3)
@@
Here, $z_{\delta}$ represents the number of neighbors of type $\delta$ (with exchange interaction $J_{\delta}$), and the summation goes over all types of neighbors. The quantity $z=1/r$ captures the effective number of neighbors, taking into account the strength ($J_{\delta}^{2}$) and number ($z_{\delta}$) of each type of interaction. Remarkably, for unfrustrated systems, the observed specific heat depends mainly on this effective number of neighbors $z$, independent of the specific lattice structure. For larger $z$ (higher coordination number), the results approach the mean-field ones, where quantum fluctuations are suppressed. Conversely, for lower $z$ (fewer neighbors), quantum fluctuations become more important, leading to a smaller specific heat below the critical temperature $T_{C}$ and a larger specific heat above $T_{C}$.
@@
Moreover, for some actual systems, the effect of the axial anisotropy $H_{A}$ becomes important, redistributing the entropy below $T_{C}$ (see Section IV).
@@
Using the above two parameters for simple lattices and the known value of $T_{C}$, we have fitted the data for the specific heat of GdNi3Ga9, GdPbBi, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. The corresponding results are presented in Section VI. Using essentially the same method, the specific heat of GdCoIn5 has been fitted in Ref. 19.
@@ II Methods
We have calculated the specific heat of a large number of magnetic structures from the numerical derivative of the internal energy as a function of temperature. The internal energy was calculated using quantum Monte Carlo (QMC) simulations with the “minbounce” algorithm (“dirloop_sse” application) from the ALPS libraries [20, 21]. Simulations were performed on systems of up to $8\times 30^{3}$ magnetic moments.
To generate the lattice configurations for the ALPS simulations, we utilized the “SUNNY” package [22] in conjunction with crystallographic information files (CIF) obtained from public repositories. Specifically, CIF files for Fe in simple cubic (SC), body-centered cubic (BCC), hexagonal close-packed (HCP), and face-centered cubic (FCC) structures, as well as C in diamond, and MgAl2Cu (for EuPdSn2) were retrieved from the Materials Project [23] database. The ErNi3Al9 CIF file was obtained from the Crystallography Open Database [24] using data from reference 25. In each case, the magnetic structure was created by replacing C, Fe, Mg, or Er with a spin moment $S=7/2$.
For all lattices, we adopted a rectangular cuboid supercell geometry as the conventional unit cell, resulting in varying numbers of sites within the supercell, with the diamond lattice requiring the largest (8 sites). Simulations were performed on systems containing $N^{3}$ supercells with periodic boundary conditions, typically with $N=10$, $20$, and $30$. For the critical temperature, results were extrapolated to the thermodynamic limit assuming a $1/N$ dependence.
The well-known sign problem [26] associated with QMC simulations limited our investigation to non-frustrated magnetic structures. At low temperatures, the number of Monte Carlo steps required for statistically reliable results increases exponentially. Consequently, the internal energy was primarily computed in a temperature range between $0.2T_{C}$ and $2T_{C}$, where $T_{C}$ is the critical temperature of the system. In a few cases, simulations were extended to $0.1T_{C}$ at low temperatures. However, the reliability of the results below $0.2T_{C}$ is limited. We used $2\times 10^{5}$ Monte Carlo steps for thermal equilibration and $6\times 10^{5}$ steps for measurements.
@@
Our theoretical results for the specific heat for all lattices used here are available in 27.
@@ III Role of the effective number of neighbors
In this Section we report our results for the specific heat and critical temperature for several simple lattices as a function of the number of neighbors that interact with a given site. For simplicity, in all lattices considered, all sites are equivalent, and we assume all $J_{\delta}=J_{1}$, where $J_{1}$ is the NN interaction, up to a certain distance $d$ and 0 for longer distances. Therefore, the effective number of neighbors $z=\Sigma_{\delta}z_{\delta}$ is just the sum of all neighbors within a sphere of radius $d$. In Section V we discuss more general cases showing that the results are not severely affected by this assumption.
@@
Most of our studies assume ferromagnetic interactions ($J_{\delta}<0$). Generally, the results for the corresponding non-frustrated antiferromagnetic structures are very similar, particularly for $z>6$, except at very low temperatures, where the specific heat dominated by magnons behaves as $T^{3/2}$ ($T^{3}$) for ferro- and antiferromagnetic structures, respectively.
@@
In Fig. 1, we show the critical temperature $T_{C}$ as a function of $z$ for the different structures considered. Calculations have been done in finite systems containing $N^{3}$ conventional unit cells, with $N=10$, 20, and 30, as explained in Section II. The extrapolated value is also indicated in the figure.
@@
As expected, $T_{C}$ increases almost linearly with $z$ as the mean-field value $T_{C}^{\text{MF}}$ (the effective magnetic field due to the interactions is proportional to $zJ_{1}$, see Section IV.1), but it is lower than $T_{C}^{\text{MF}}$, particularly for low $z$ due to the effect of quantum fluctuations, which is larger for low $z$.
For large enough $z$ one expects that $T_{C}$ and $T_{C}^{\text{MF}}$ coincide. However, for large $z$ in a finite system with periodic boundary conditions, the effective coordination is smaller when the range of the interactions $d$ becomes of the order of the system size, because neighbors at different distances coincide, and it is not possible to reach this limit with our method.
@@
Figure 2: (Color online) Specific heat over temperature as a function of temperature for several ferromagnetic structures with different effective number of neighbors $z$. Top: $z\leq 12$. Bottom: $z\geq 12$. There is an almost perfect superposition between HCP and FCC for $z=12$, among SC, FCC and HCP for $z=18$, and among SC, BCC and the EuPdSn2 structure for $z=26$. Therefore, only one of the curves for each of these three sets is shown. LHc(4) corresponds to the layered honeycomb structure with $z=4$. Bottom inset shows the specific heat for FCC(12) lattices with sizes $N=10$ (black line), $20$ (red), and $30$ (green) in the range $1.50.5<T/T_{C}<$ 1.5.
@@
In Fig. 2 we show the results for the specific heat as a function of temperature for a large number of ferromagnetic structures with different numbers of magnetic neighbors $z$. Also shown is the mean-field result which corresponds to the limit $z\rightarrow\infty$. This result shows larger $C$ for $T<T_{C}$, while $C=0$ for $T>T_{C}$. In contrast, the specific heat for a honeycomb layered system with $z=4$ (three nearest neighbors within the plane of the honeycomb layer and one above or below the layer in alternating sites) shows the smallest specific heat below $T_{C}$ and the largest one above $T_{C}$. Note that the diamond structure, which is completely different from the previous one but also has $z=4$, has a very similar specific heat. The low-temperature shoulder is a Schottky-like effect arising from the internal field [19, 15]. The lower panel inset shows that finite size effects are only appreciable close to $T_{C}$ as exemplified in the case of the FCC(12) lattice. For this reason, we show specific heat curves from QMC corresponding to $N=20$.
@@
Increasing $z$ from 4 to 6, the specific heat of the simple cubic structure with six NN displays the expected trend of increasing (decreasing) $C$ below (above) $T_{C}$. Increasing $z$ up to 14 using the BCC structure with 6 NN and 8 next-nearest neighbors (NNN), the same trend is observed in general. However, the specific heat of the honeycomb lattice with $z=10$ (four links as described before and adding 6 second neighbors in the plane) lies slightly below the cubic BCC with $z=8$ nearest neighbors for $T\sim 0.5T_{C}$. In addition, somewhat unexpectedly, the result for the simple cubic structure including 6 NN and 8 third NN, shows that $C$ at $T\sim 0.5T_{C}$ is greater than the corresponding one of the above-mentioned BCC structure with the same $z=14$, and also greater than the specific heat of one structure with $z=16$ and two structures with $z=18$. Nevertheless, except for this case and the above-mentioned inversion between $z=8$ and 10, the specific heat is very similar for different structures with the same $z$ and increases with increasing $z$ at intermediate temperature $T\sim 0.5T_{C}$.